# MT2KD: Towards A General-Purpose Encoder for Speech, Speaker, and Audio Events

### v1

@@ V-C Consistency of Data Augmentation
The effect of applying different augmentations to the student model input data was investigated here and the results are shown in Table V. Different augmentations applied to the input of student model input when performing KD give rise to different effects on different tasks. Applying SpecAug [43] to the input features of the student model lead to weaker KD performance on ASR and SV, but slightly improves AT performance. Mixing input features with Musan [44] for noise augmentation leads to better SV performance, but undermines the KD result on ASR and AT. Since the teacher targets were obtained using the un-augmented inputs, not applying augmentations to the student model input ensures consistency during KD training of the student model. Note that consistent data augmentation is possible in on-line KD, but the training efficiency is greatly affected due to the extra computation cost of teacher models. To minimise the KD performance reduction on all three tasks, no augmentation was performed to the student model input in the following multi-teacher KD experiments.
@@ V-D1 ASR + AT
Multi-task KD training experiments for ASR and AT were performed and results are shown in Table VI. As reported in Whisper-AT, the last few layers of the Whisper model encodes the sound event information, suggesting that ASR and AT are not affecting each other. Since Whisper is selected as the ASR teacher, the embeddings of the last layer of the student encoder was used as the input to the AT module. As shown in Table VI, performing multi-task KD on ASR and AT leads to only minor degradation on each task.
@@ V-E Fine-tuning
Although KD pre-training is performed using unlabelled data, the student model also possesses the capability of performing AT and SV due to the KD loss function for these two tasks after KD. However, the whole system still needs to be fine-tuned with ASR data. The main goal during the fine-tuning is to give the model good ASR performance while having minimal impact on AT and SV. Experiments were carried out on with different initialisations. For easier reference, the model names of the KD pre-trained models are listed in Table X.
TABLE X: The names of the KD pre-trained models.
Pre-training Data Model Name
@@
Since parameters involving SV only make up a small proportion of the whole encoder, one option would be freezing the parameters involving SV and thus only the rest of the model for other tasks are fine-tuned. Also, it is shown in Table IV that using KD achieves better performance on SV and AT than supervised training, the KD loss can be still be adopted when fine-tuning the whole model on ASR tasks. The results of different fine-tuning strategies for SV are shown in Table XI. All models were fine-tuned with LS-960 for ASR, AS-2M for AT and Vox1,2 for SV (if SV parameters are not frozen).
@@
If the whole model is fine-tuned without using KD as an auxiliary loss, the model already outperforms the multi-task baseline model trained from scratch (last row in Table III) by a large margin on all three tasks, validating the effectiveness of multi-teacher KD pre-training. If the KD loss is adopted as an auxiliary loss during fine-tuning, the mAP and EER are improved at the cost of big WER increase. Note that the mAP and EER are worse than the initialisation model. This aligns with the observation (see Table III) that multi-task learning leads to weaker performance on individual tasks. In particular, it is found that SV performance drops significantly compared to the initialisation model. To address this problem, another fine-tuning strategy is explored, where the SV-related parameters are frozen during the fine-tuning process to maintain the same SV performance as the initialisation model after fine-tuning. As shown in the final two rows in Table XI, ASR performance of the partially fine-tuned model is slightly worse than the fully fine-tuned model, but the overall performance on the three tasks is better. Our final system uses the best pre-trained model (ST3) as an initialisation, and is partially fine-tuned with the ASR loss and KD loss for AT. The performance of the final system is shown in the last row of Table XI. It achieved WERs of 2.35/5.63, mAP of 45.9 and EER of 1.13. Compared to the best-performing single-task teacher models, the multi-task student model yields competitive performance on all three tasks while requiring for fewer parameters and reduced computation.
@@ VI Conclusions
This paper presents a novel two-stage multi-task learning framework to build a general-purpose speech and audio encoder that jointly performs three fundamental tasks: ASR, AT and SV. In the first stage, multi-teacher KD is applied to align the feature space of a student model to three teacher models specialised for each task using unlabelled data. In the second stage, multi-task supervised fine-tuning is performed by initialisating the model from the first stage. Experiments show that the proposed framework outperforms a multi-task model trained from scratch with supervised data by a large margin. The final multi-task model achieives WERs (%) of 2.35/5.63 on LibriSpeech test sets, mAP of 45.9% on AudioSet and EER (%) of 1.13 on VoxCeleb1 test set, which is close to the best-performing single-task models while requiring for fewer parameters and less computation.

### v3

@@ V-C Consistency of Data Augmentation
The effect of applying different augmentations to the student model input data was investigated here and the results are shown in Table V. Different augmentations applied to the input of student model input when performing KD give rise to different effects on different tasks. Applying SpecAug [43] to the input features of the student model lead to weaker KD performance on ASR and SV, but slightly improves AT performance. Mixing input features with Musan [44] for noise augmentation leads to better SV performance, but undermines the KD result on ASR and AT. Since the teacher targets were obtained using the un-augmented inputs, not applying augmentations to the student model input ensures consistency during KD training of the student model. Note that consistent data augmentation is possible in on-line KD, but the training efficiency is greatly affected due to the extra computation cost of teacher models. To minimise the KD performance reduction on all three tasks, no augmentation was performed to the student model input in the rest multi-teacher KD experiments.
@@ V-D1 ASR + AT
Multi-task KD training experiments for ASR and AT were performed and results are shown in Table VI. As reported in Whisper-AT, the last few layers of the Whisper model encodes the sound event information, suggesting that ASR and AT are not affecting each other. Since Whisper is selected as the ASR teacher, the embeddings of the last layer of the student encoder were used as the input to the AT module. As shown in Table VI, performing multi-task KD on ASR and AT leads to only a minor degradation on each task.
@@ V-E Fine-tuning
Although KD pre-training is performed using unlabelled data, the student model also possesses the capability of performing AT and SV due to the KD loss function for these two tasks after KD. However, the whole system still needs to be fine-tuned with ASR data. The main goal during the fine-tuning is to give the model good ASR performance while having minimal impact on AT and SV. Experiments were carried out on with different initialisations. For easier reference, the model IDs of the KD pre-trained models are listed in Table X.
TABLE X: The IDs of the KD pre-trained models.
Pre-training Data Model ID
@@
Since parameters involving SV only make up a small proportion of the whole encoder, one option would be freezing the parameters involving SV and therefore only the rest of the model for other tasks are fine-tuned. Also, it is shown in Table IV that using KD achieves better performance on SV and AT than supervised training, the KD loss can be still be adopted when fine-tuning the whole model on ASR tasks. The results of different fine-tuning strategies for SV are shown in Table XI. All models were fine-tuned with LS-960 for ASR, AS-2M for AT and Vox1,2 for SV (if SV parameters are not frozen).
@@
If the whole model is fine-tuned without using KD as an auxiliary loss, the model already outperforms the multi-task baseline model trained from scratch (last row in Table III) by a large margin on all three tasks, validating the effectiveness of multi-teacher KD pre-training. If the KD loss is adopted as an auxiliary loss during fine-tuning, the mAP and EER are improved at the cost of a large WER increase. Note that the mAP and EER are worse than for the initialisation model. This aligns with the observation (see Table III) that multi-task learning leads to weaker performance on individual tasks. In particular, it is found that SV performance drops significantly compared to the initialisation model. To address this problem, another fine-tuning strategy is explored, where the SV-related parameters are frozen during the fine-tuning process to maintain the same SV performance as the initialisation model after fine-tuning. As shown in the final two rows in Table XI, ASR performance of the partially fine-tuned model is slightly worse than the fully fine-tuned model, but the overall performance on the three tasks is better. Our final system uses the best pre-trained model (ST3) as an initialisation, and is partially fine-tuned with the ASR loss and KD loss for AT. The performance of the final system is shown in the last row of Table XI. It achieved WERs of 2.35/5.63, mAP of 45.9 and EER of 1.13. Compared to the best-performing single-task teacher models, the multi-task student model yields competitive performance on all three tasks while requiring far fewer parameters and reduced computation.
@@ VI Conclusions
This paper presents a novel two-stage multi-task learning framework to build a general-purpose speech and audio encoder that jointly performs three fundamental tasks: ASR, AT and SV. In the first stage, multi-teacher KD is applied to align the feature space of a student model to three teacher models specialised for each task using unlabelled data. In the second stage, multi-task supervised fine-tuning is performed by initialising the model from the first stage. Experiments show that the proposed framework outperforms a multi-task model trained from scratch with supervised data by a large margin. The final multi-task model achieives WERs (%) of 2.35/5.63 on the LibriSpeech test sets, mAP of 45.9% on AudioSet, and EER (%) of 1.13 on the VoxCeleb1 test set, which is close to the best-performing single-task models while requiring far fewer parameters and less computation.